OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 1 −2 −5 −8

Mehek (mehek14):

do you know how t find the slope?

OpenStudy (anonymous):

not really

Mehek (mehek14):

for the fist y value, are you sure it's 1?

OpenStudy (campbell_st):

as x goes up by 1's what does y decrease by..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by 3 i think

OpenStudy (campbell_st):

well its actually -3 thats the slope... each y value changes by -3 so the equation is y = -3x + b to find the value of b, just substitute any of the points say x = 1 and y = 1 so 1 = -3*1 + b find b

OpenStudy (anonymous):

wait i don't get it

OpenStudy (anonymous):

help!! ;)

OpenStudy (campbell_st):

ok, the change in the y's as x change by 1's is the slope...

OpenStudy (lizz123):

The equation he used was \[m=\frac{ y1-y2 }{ x2-x1 }\]

OpenStudy (campbell_st):

so the slope intercept form of the equation is y = mx + b m = sloep and b = y- intercept you found the slope m = -3 so the equation is y = -3x + b you have a table of points (1, 1) is a point on the line so substitute the values x = 1 and y = 1 into the equation to find b so you need to solve \[1 = -3 \times 1 + b\] what value do you think b has...?

OpenStudy (lizz123):

Choose two points doesn't matter just choose two

OpenStudy (anonymous):

Slope = the change in the rise ( change in Y) divided by the change in the run (change in X). For this problem, Lizz123 gave you the correct formula. Just make sure that you use the correct numerators (Y's) and denominators (X's)

OpenStudy (anonymous):

so i think b= 1 or 2 i think

OpenStudy (lizz123):

No, there is a pattern in the y-axis.

OpenStudy (campbell_st):

close here is the table and you can use that 

OpenStudy (anonymous):

the pattern is it decreasing by -3

OpenStudy (lizz123):

Yes.

OpenStudy (anonymous):

When I teach about slope I don't use b. Just use two points on the graph and do the simple formula that Lizz123 gave you and it is a breze.

OpenStudy (campbell_st):

so ? - 3 = 1 what is ?

OpenStudy (lizz123):

So, now your answer is what?

OpenStudy (lizz123):

@tigergirl0000

OpenStudy (campbell_st):

Question @fribby.... why do you want to fill people's heads with formula... when understanding patterns achieves the same result -------------

OpenStudy (anonymous):

the answer is y= -3x + 1 or 0 idk

OpenStudy (lizz123):

no

OpenStudy (campbell_st):

ok.... its close.... its just finding the value of b... called the y- intercept.... can you tell me the missing number in the sum ___- 3 = 1

OpenStudy (anonymous):

2?

OpenStudy (lizz123):

@tigergirl0000 you said b=-3 so using what you said: y=-3-/+b

OpenStudy (campbell_st):

2 - 3 = -1 can you try again..

OpenStudy (lizz123):

What is b?

OpenStudy (campbell_st):

ok.... does this make sense 4 - 3 = 1....?

OpenStudy (anonymous):

yea that makes sense

OpenStudy (anonymous):

does b= -3

OpenStudy (lizz123):

yes

OpenStudy (lizz123):

write the equation y=-3-?

OpenStudy (campbell_st):

 so the missing value in the table is 4... the point (0,b) is the y- intercept and in the table b = 4 so the equation is y = -3x + 4